OpenStudy (anonymous):

Prove whether f : Z × Z → Z is one-to-one AND onto if f (m, n) = m + n + 1. Thanks!

OpenStudy (amoodarya):

are u sure that is one-to-one ? f(5,3)=5+3+1=9 f(4,4)=4+4+1=9 ??

OpenStudy (amoodarya):

but onto is ok because for example f(-1,x)=-1+x+1=x x belongs to Z so , is onto function

OpenStudy (zzr0ck3r):

it is for sure NOT one-to-one in \(\mathbb{R}\). Example \(f(4,6)=f(3,7)\) but \((4,6)\ne(3,7)\)

OpenStudy (zzr0ck3r):

disregard the "in \(\mathbb{R}\)"

OpenStudy (zzr0ck3r):

for onto Suppose \(z\in \mathbb{Z} \), then \(z\) is either even or odd if \(z\) is even let \(x=\frac{z}{2}\) and \(y=\frac{z}{2}-1\) then \(f(x,y)=z\) if \(z\) is odd, let \(x=y=\frac{z-1}{2}\) then \(f(x,y)=z\).

OpenStudy (anonymous):

oh yeah lol we had to prove whether or not it is one on one. I proved it is onto, but not one to one